An 77.0 kg spacewalking astronaut pushes off a 645 kg satellite, exerting a 85.0 N force for the 0.500 s it takes him to straighten his arms.

How far apart are the astronaut and the satellite after 1.30 min ?
Answer #1

Sol::
Given
Msat= 645 kg
Mas=77 kg
F= 85 N
t=0.5 s
Now using conservation of linear momentum,

F_action = (M_sat) Δv/Δt

85 = (645)(v - 0)/0.5

v = 0.0658 m/s

F_reaction = -(M_ast) Δv/Δt

85 = -(77)(V - 0)/0.5

V = -0.552 m/s

relative velocity satellite with respect to astronaut is,

ΔV = v - V

ΔV = 0.0658 - (-0.552)

ΔV = 0.6178 m/s

s = ΔV t

s = 0.6178* (1.3 * 60)

s = 48.188 m
s =48.2 m


Hope this helps and will get RATED

THANK YOU